# UMGAD: Unsupervised Multiplex Graph Anomaly Detection

### v2

@@ VI Conclusion
In this work, we propose UMGAD, a novel framework tailored for UGAD tasks. To comprehensively capture diverse anomaly signals, UMGAD employs a collaborative network that integrates graph autoencoders with diverse masking strategies across multiple relation types. These strategies operate effectively in both original and augmented views, enabling the model to extract meaningful patterns while filtering out redundant information. Leveraging attribute and structure reconstruction in the augmented view, followed by anomaly aggregation through dual-view contrastive learning, UMGAD achieves robust and adaptive detection. Extensive experiments on datasets with synthetic and real-world anomalies demonstrate that UMGAD outperforms state-of-the-art methods in unsupervised anomaly detection, excelling in both performance and efficiency. Ablation studies further validate the contribution of each component. In future work, we aim to explore simpler yet highly efficient models to push the boundaries of UGAD tasks.

### v4

@@ VI Conclusion
In this paper, we propose UMGAD, an unsupervised framework for anomaly detection on multiplex heterogeneous graphs. It combines different relation-aware masking mechanisms with graph autoencoders to capture informative signals while reducing redundancy. Attribute and structure reconstruction on both original and augmented views, followed by dual-view contrastive learning, ensures robust anomaly detection. Extensive experiments on datasets with injected and real-world anomalies show that UMGAD outperforms state-of-the-art methods in accuracy and efficiency.
acknowledge
This work is partially supported by the National Natural Science Foundation of China under grant Nos. 62176243 and 62472263, the National Science and Technology Major Project under grant No. 2024ZD0607700, the China Postdoctoral Science Foundation (No. 2024T170868), the Postdoctoral Fellowship Program of CPSF (No. GZC20232500), and the Shandong Postdoctoral Science Foundation (No. SDCX-ZG-202400317).